a copper wire has a circular cross section with a radius of 1.25 mm. a) if the wire carrries a current of 3.70 A, FIND THE DRIFT SPEED OF THE ELECTRONS in this wire. b)all other things being equal what happens to the drift speed in wires made of metal having a larger number of conduction electrons per atom than copper ? Explain
1
Expert's answer
2020-04-09T09:24:01-0400

a)


"v=\frac{I}{neA}=\frac{I}{\pi r^2ne}"

"v=\frac{3.7}{\pi(0.00125)^2(8.46\cdot10^{28})(1.6\cdot10^{-19})}=5.57\cdot10^{-5}\frac{m}{s}"

b)  The drift speed will decrease.


"v'=\frac{I}{n'eA}=\frac{n}{n'}v<v"